# Spatial Conformal Inference through Localized Quantile Regression

### v1

@@ Abstract
Reliable uncertainty quantification at unobserved spatial locations, especially in the presence of complex and heterogeneous datasets, remains a core challenge in spatial statistics. Traditional approaches like Kriging rely heavily on assumptions such as normality, which often break down in large-scale, diverse datasets, leading to unreliable prediction intervals. While machine learning methods have emerged as powerful alternatives, they primarily focus on point predictions and provide limited mechanisms for uncertainty quantification. Conformal prediction, a distribution-free framework, offers valid prediction intervals without relying on parametric assumptions. However, existing conformal prediction methods are either not tailored for spatial settings, or existing ones for spatial data have relied on rather restrictive i.i.d. assumptions. In this paper, we propose Localized Spatial Conformal Prediction (LSCP), a conformal prediction method designed specifically for spatial data. LSCP leverages localized quantile regression to construct prediction intervals. Instead of i.i.d. assumptions, our theoretical analysis builds on weaker conditions of stationarity and spatial mixing, which is natural for spatial data, providing finite-sample bounds on the conditional coverage gap and establishing asymptotic guarantees for conditional coverage. We present experiments on both synthetic and real-world datasets to demonstrate that LSCP achieves accurate coverage with significantly tighter and more consistent prediction intervals across the spatial domain compared to existing methods.
Keywords: conformal prediction, uncertainty quantification, spatial statistics, mixing
@@ 3 Method
In this section, we will first introduce the background of conformal prediction and our proposed Local Spatial Conformal Prediction (LSCP) method. Then we will discuss the benefits of LSCP compared to several alternatives.
@@ 3.2 Proposed method: Local spatial conformal prediction (LSCP)
The empirical quantile places an equal weight on all the points, which may not fully capture the dependence structure. Instead, we apply a quantile regression estimator $\hat{Q}_{n}$ on the residuals $\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})}$. The estimator, $\hat{Q}_{n}(\alpha)$, predicts the $\alpha$-quantile of the residual $\hat{\varepsilon}(s_{n+1})$ given the values of its neighbors. For computational efficiency, we use Quantile Random Forests from (Meinshausen & Ridgeway, 2006), although other quantile regression techniques could also be applied. The resulting prediction interval is:
$\hat{C}_{n}(X(s_{n+1}))=\left[\hat{f}(X(s_{n+1}))+\hat{Q}_{n}(\beta^{*}),\;\hat{f}(X(s_{n+1}))+\hat{Q}_{n}(1-\alpha+\beta^{*})\right],$ (1)
where $\beta^{*}=\operatorname{argmin}_{\beta\in[0,\alpha]}(\hat{Q}_{n}(1-\alpha+\beta)-\hat{Q}_{n}(\beta))$. Here $\beta$ is optimized to find the tightest interval.
@@
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i\in N(s_{n+1})}\omega_{i}\mathbf{1}\{\tilde{Y}(s_{i})\leq e\}\leq p\},$
@@
Figure 2: Illustration of Localized Spatial Conformal Prediction. Left: $k$-nearest neighbors with weights are used to predict the test point while providing uncertainty quantification. Right: The $k$-nearest neighborhood adapts to both dense and sparse observation regions. Red points represent test locations, with dashed circles marking neighborhoods of the 5-nearest neighbors. Darker grey shades indicate regions of lower uncertainty. LSCP leverages these neighboring points to construct prediction intervals at the test locations.
@@ 3.3 Comparison with related methods
3.3.1 Global Spatial Conformal Prediction (GSCP)
GSCP, introduced in (Mao et al., 2024), applies equal weighting to all non-conformity scores across the calibration dataset, where the non-conformity score is defined as
@@ 3.3.1 Global Spatial Conformal Prediction (GSCP)
7:Set $\tilde{Y}(s_{i})=\hat{\varepsilon}(s_{i})$ and $\tilde{X}(s_{i})=(\hat{\varepsilon}(s_{j_{1}}),\dots,\hat{\varepsilon}(s_{j_{|N(s_{i})|}}))$, where $s_{j}\in N(s_{i})$.
8:Fit quantile regression $\hat{Q}_{n}$ with all pairs $(\tilde{X}(s_{i}),\tilde{Y}(s_{i}))$ in the calibration data.
9:Obtain the prediction interval $\hat{C}_{n}(X(s_{n+1}))$.
3.3.2 Smoothed Local Spatial Conformal Prediction (SLSCP)
SLSCP, introduced in (Mao et al., 2024), improves upon previous GSCP by utilizing only nearby data to construct prediction intervals. The non-conformity scores are defined as in Equation 2, but the quantile is estimated locally:
@@ 3.3.2 Smoothed Local Spatial Conformal Prediction (SLSCP)
3.3.3 Localized Conformal Prediction (LCP)
LCP, introduced in (Guan, 2023), provides a general framework for localized conformal prediction rather than the spatial setting only. The method combines GSCP and SLSCP in quantile estimation:
@@ 3.3.3 Localized Conformal Prediction (LCP)
LSCP (Ours) SLSCP ((Mao et al., 2024)) LCP ((Guan, 2023))
@@
Weights Learned by quantile regression
Specifed kernel function on
location $s$
Specifed kernel function on
feature $X(s)$
Assumption on data distribution
Stationary and spatial mixing
noise $\varepsilon(s)$
Locally i.i.d. noise $\varepsilon(s)$ Globally i.i.d. data $(X(s),Y(s))$
Assumption on data model Additive model
Continuous mapping from
$L_{2}$-continuous spatial process
and noise process
No assumption
@@ 4.1 Setting
$Y(s)=f(X(s))+\varepsilon(s),$ (5)
where $f$ is an unknown function and $\varepsilon(s)$ represents the noise process, whose marginal distribution is not necessarily Gaussian. Given a pretrained prediction model $\hat{f}$, we can compute the non-conformity scores
@@ 4.2 Preliminary
For simplifying the exposition, we further assume that there exist a nonincreasing function $\alpha_{1}(\cdot)$ with $\lim_{a\rightarrow\infty}\alpha_{1}(a)=0$ and a nondecreasing function $g(\cdot)$ such that the strong-mixing coefficients $\alpha(a,b)$ satisfies the inequality
@@ 4.3 Assumptions
With the data $(X(s_{1}),Y(s_{1})),\cdots,(X(s_{n}),Y(s_{n}))$, we would like to construct a prediction region for $Y(s_{n+1})$ where only the feature $X(s_{n+1})$ is known. Define $\omega_{n}(x)$ as the weight function and $\widetilde{\omega}_{n}(X(s_{i}))$ to be the normalized weight over samples $s_{1},\cdots,s_{n}$ so that $\sum_{i=1}^{n}\widetilde{\omega}_{n}(X(s_{i}))=1$. The weight function can be some function that measures the similarity between $X(s)$ and $X(s_{n+1})$.
@@
$\widetilde{\omega}_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}}),$ (7)
for all $i$. In other word, $M_{n}=\max_{1\leq i\leq n}\widetilde{\omega}_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}})$.
The requirement assumes that the normalized weights decay at a rate faster than $n^{-\frac{1}{2}}$. As we can see, $\widetilde{\omega}_{n}=\frac{1}{n}$ is a special case that satisfies this condition. The condition is also weaker than the requirement of $\widetilde{\omega}_{n}=O(\frac{1}{n})$ in a related study (Xu et al., 2024). Besides, the assumption can also be inferred from that of SLSCP (Mao et al., 2024) where a GBF kernel with infinitely close data leads to uniform weights.
@@
$\sum_{i=1}^{n}\|\hat{\varepsilon}(s_{i})-\varepsilon(s_{i})\|^{2}\leq\frac{\delta_{n}^{2}}{M_{n}},~{}~{}\|\hat{\varepsilon}(s_{n+1})-\varepsilon(s_{n+1})\|\leq\delta_{n}.$ (8)
@@ 4.4 Coverage guarantee
Lemma 4.4.
(Distance between the empirical CDF of $\varepsilon$ and $\hat{\varepsilon}$) Under Assumption 4.1 and 4.2,
@@
Lemma 4.5.
(Convergence of empicial CDF of $\varepsilon$) Under Assumptions 4.1-4.3, with probability higher than $1-(2+M+2\sqrt{M}g(b))n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}$,
@@
where $\widetilde{F}_{n+1}(y)=\sum_{i=1}^{n}\widetilde{\omega}_{n}(X(s_{i}))\mathbf{1}(\epsilon(s_{i})\leq y)$.
Our main theorem is the following one, which establish the asymptotic convergence of conditional coverage using the previous lemma.
Theorem 4.6.
(Conditional Coverage) Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
@@
Corollary 4.7.
(Marginal Coverage) Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
@@
5 Experiments
In this section, we compare our proposed LSCP method against four baseline approaches. The first method, EnbPI ((Xu & Xie, 2021)), is designed for time series data and applies equal weight to the most recent observations. The second method, GSCP ((Mao et al., 2024)), uses the entire dataset for prediction, applying equal weights to all data points. The third method, SLSCP ((Mao et al., 2024)), leverages k-nearest neighbors for uncertainty quantification, assigning different weights to each point based on the spatial distance between the test data and the calibration data. The fourth method, LCP ((Guan, 2023)), utilizes all data points for prediction, weighting them according to the similarity between the features of test data and calibration data. Both SLSCP and LCP use the Gaussian kernel as a similarity measure.
5.1 Synthetic data experiments
We begin by comparing our method, LSCP, with baseline methods across several simulated scenarios. For simplicity, we assume that the locations $s$ are uniformly sampled from the unit grid $[0,1]\times[0,1]$. The mean-zero stationary Gaussian process $X(s)$ is defined with a Matérn covariance function, having variance $\sigma^{2}=1$, range $\phi=0.1$, and smoothness $\kappa=0.7$. The scenarios are as follows:
  1. 1.
    $Y(s)=X(s)+\epsilon(s)$.
  2. 2.
    $Y(s)=X(s)|\epsilon(s)|$.
  3. 3.
    $Y(s)=X(s)+\sin(\|s\|_{2})\epsilon(s)$.
These scenarios include nonlinear and non-stationary settings that extend beyond the assumptions made in our theoretical framework. However, the empirical results demonstrate that our proposed LSCP method consistently outperforms other methods across all scenarios.
In Table 2, LSCP achieves the target coverage of $90\%$ across all cases while also maintaining a significantly narrower prediction interval. Figure 3 displays the interval width over the spatial domain for each scenario and the true residuals over the space. GSCP is omitted from these plots as it produces a uniform interval width across the space. A larger interval width indicates higher model uncertainty. These figures reveal that LSCP adapts more locally, while the other methods display smoother, more uniform patterns. Both EnbPI and SLSCP rely on k-nearest neighbors for constructing prediction intervals, differing primarily in their choice of weights. As bandwidth increases, the performance of SLSCP approximates that of EnbPI, since SLSCP weights approaching uniformity. A similar pattern is observed between LCP and GSCP.
To further assess the methods, we divide the unit grid into $10\times 10$ smaller areas, computing the coverage and average interval width for each method within each area. The resulting distributions are shown in Figure 5. LSCP and GSCP maintain an average coverage above the target $90\%$, with LSCP showing the most concentrated distribution, indicating stability and consistency across areas. In the average width plot, LSCP shows the smallest width while being the most consistent. On the other hand, LCP displays a similar performance compared to GSCP, and a major reason is that they both use the whole calibration dataset for prediction. When the bandwidth of the kernel grows larger, the weights in LCP converge to $\frac{1}{n}$ in GSCP. Overall, our proposed LSCP not only provides tighter valid regions but also demonstrates the highest consistency.
@@ 5.1 Synthetic data experiments
Figure 3: The heatmaps illustrate the width of the prediction intervals for each method across the three scenarios.
Table 2: Simulation: The table presents a comparison of the coverage and prediction interval width for five methods across three scenarios. LSCP consistently achieves the target coverage of $90\%$ in all cases, while also maintaining the smallest prediction interval width. The number in the bracket shows the bandwidth selected by $5$-fold cross-validation. When the bandwidth is large, the performance of LCP is similar to GSCP, and the performance of SLSCP is close to EnbPI.
@@
SLSCP 90.6% ($0.03$) 1.64 88.2% ($0.03$) 0.73 89.1% ($0.03$) 1.07
LCP 89.6% ($0.5$) 1.92 94.1% ($0.02$) 0.78 91.8% ($0.01$) 1.43
(a) New Mexico
(b) Georgia
Figure 4: Mobile signal score from Ookla dataset. The plots show the distribution of data accross the state.
5.2 Real data experiments
In the real data experiments, we utilize the mobile network measurement data from the Ookla public dataset, which records user-reported statewide mobile internet performance. Our analysis focuses on the New Mexico and Georgia dataset. Figure 4 illustrates the spatial distribution of data across the two states. As we can see, the data distributes unevenly across urban and suburb areas. We employ kernel regression as the prediction method $\hat{f}$, and compare the performance of the conformal methods. The New Mexico dataset contains $24983$ observations in total and the Georgia dataset has $28587$ data points.
As summarized in Table 3, our proposed LSCP method demonstrates a clear advantage over competing methods, offering both significantly narrower prediction intervals and superior coverage. Similar to the synthetic experiments, we divide each state into $10\times 10$ uniform grids and calculate the coverage and average interval width for the test data within each grid. These metrics allow us to examine the spatial consistency of the methods. The violin plots in Figure 5 visually represent the distribution of coverage and interval width across the grids. It is important to note that the averages presented in the violin plots differ from those in Table 3. While the table reports averages over all test points, the violin plots reflect the average of grid areas. Consequently, the violin plots offer a different perspective on how the methods perform spatially, highlighting whether they behave uniformly across the entire state. The coverages reported in the table are generally higher because these methods perform well in the urban areas, where data is more concentrated.
The results reveal that LSCP consistently achieves high coverage while maintaining a narrow and spatially uniform interval width, outperforming alternative approaches. This is particularly evident in the violin plots, where LSCP exhibits one of the highest and most consistent coverage rates and the most uniform interval widths across the state. This underscores the robustness and adaptability of LSCP to non-uniform data distributions, offering reliable and geographically consistent prediction intervals. These findings align with those from our synthetic study, further validating the effectiveness of LSCP in diverse and challenging data settings.
Scenario 1
Scenario 2
Scenario 3
New Mexico
Georgia
(a) Coverage
(b) Width
Figure 5: The violin plots on the left show the distribution of coverage across different areas, while the plots on the right show the distribution of width. Each row represents a different scenario.
Table 3: Real data: The table presents a comparison of the coverage and prediction interval width for five methods on mobile signal data. LSCP reaches target coverage with the smallest prediction interval width. Since the bandwidth chosen by CV is large, the performance of LCP is similar to GSCP, and the performance of SLSCP is close to EnbPI.
@@ 5.2 Real data experiments
SLSCP (bd=0.2) 89.8% 276.8 89.6% 167.8
LCP (bd=0.1) 89.7% 266.4 89.3% 166.7
6 Discussion

### v2

@@ Abstract
Reliable uncertainty quantification at unobserved spatial locations, particularly for complex and heterogeneous datasets, is a key challenge in spatial statistics. Traditional methods like Kriging rely on assumptions such as normality, which often fail in large-scale datasets, leading to unreliable intervals. Machine learning methods focus on point predictions but lack robust uncertainty quantification. Conformal prediction provides distribution-free prediction intervals but often assumes i.i.d. data, which is unrealistic for spatial settings. We propose Localized Spatial Conformal Prediction (LSCP), designed specifically for spatial data. LSCP uses localized quantile regression and avoids i.i.d. assumptions, instead relying on stationarity and spatial mixing, with finite-sample and asymptotic conditional coverage guarantees. Experiments on synthetic and real-world data show that LSCP achieves accurate coverage with tighter, more consistent prediction intervals than existing methods.
@@ 3.2 Proposed method: Local spatial conformal prediction (LSCP)
The empirical quantile places an equal weight on all the points, which may not fully capture the dependence structure. Instead, we apply a quantile regression estimator $\widehat{Q}_{n}$ on the residuals $\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})}$. The estimator, $\widehat{Q}_{n}(\alpha)$, predicts the $\alpha$-quantile of the residual $\hat{\varepsilon}(s_{n+1})$ given the values of its neighbors. For computational efficiency, we use Quantile Random Forests from (Meinshausen & Ridgeway, 2006), although other quantile regression techniques could also be applied. The resulting prediction interval is:
$\hat{C}_{n}(X(s_{n+1}))=[\hat{f}(X(s_{n+1}))+\widehat{Q}_{n}(\beta^{*}),\;\hat{f}(X(s_{n+1}))+\widehat{Q}_{n}(1-\alpha+\beta^{*})],$
where $\beta^{*}=\operatorname{argmin}_{\beta\in[0,\alpha]}(\widehat{Q}_{n}(1-\alpha+\beta)-\widehat{Q}_{n}(\beta))$. Here $\beta$ is optimized to find the tightest interval.
@@
$\widehat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i\in N(s_{n+1})}\omega_{i}\mathbf{1}\{\tilde{Y}(s_{i})\leq e\}\leq p\},$
@@
Figure 2: Illustration of Localized Spatial Conformal Prediction (LSCP). Left: $k$-nearest neighbors with weights predict test points and quantify uncertainty. Right: Neighborhoods adapt to dense and sparse regions. Red points mark test locations, dashed circles show 5-nearest neighbors, and darker grey indicates lower uncertainty. LSCP uses neighbors to construct prediction intervals.
@@ 3.3 Comparison with related methods
3.3.1 Global Spatial Conformal Prediction
Global Spatial Conformal Prediction (GSCP), introduced in (Mao et al., 2024), applies equal weighting to all non-conformity scores across the calibration dataset, where the non-conformity score is defined as
@@ 3.3.1 Global Spatial Conformal Prediction
7:Set $\tilde{Y}(s_{i})=\hat{\varepsilon}(s_{i})$ and
8:$\tilde{X}(s_{i})=(\hat{\varepsilon}(s_{j_{1}}),\dots,\hat{\varepsilon}(s_{j_{|N(s_{i})|}}))$, where $s_{j}\in N(s_{i})$.
9:Fit quantile regression $\widehat{Q}_{n}$ with all pairs $(\tilde{X}(s_{i}),\tilde{Y}(s_{i}))$ in the calibration data.
10:Obtain the prediction interval $\hat{C}_{n}(X(s_{n+1}))$.
3.3.2 Smoothed Local Spatial CP
Smoothed Local Spatial Conformal Prediction (SLSCP) (Mao et al., 2024) improves upon previous GSCP by utilizing only nearby data to construct prediction intervals. The non-conformity scores are defined as in Equation 1, but the quantile is estimated locally:
@@ 3.3.2 Smoothed Local Spatial CP
3.3.3 Localized Conformal Prediction
Localized Conformal Prediction (LCP), introduced in (Guan, 2023), provides a general framework for localized conformal prediction rather than the spatial setting only. The method combines GSCP and SLSCP in quantile estimation:
@@ 3.3.3 Localized Conformal Prediction
LSCP (Ours) SLSCP (Mao et al., 2024) LCP (Guan, 2023)
@@
Weights learned by quantile regression parametric kernel for location $s$ parametric kernel for feature $X(s)$
Distributional assumption stationary and spatial mixing noise $\varepsilon(s)$ locally i.i.d. noise $\varepsilon(s)$ globally i.i.d. data $(X(s),Y(s))$
Data model additive noise continuous mapping from $L_{2}$-continuous spatial and noise processes no assumption
@@ 4.1 Setting
$Y(s)=f(X(s))+\varepsilon(s),$
where $f$ is an unknown function and $\varepsilon(s)$ represents the noise process, whose marginal distribution is not necessarily Gaussian. Given a pre-trained prediction model $\hat{f}$, we can compute the non-conformity scores
@@ 4.2 Preliminary
To simplify the exposition, we further assume that there exists a nonincreasing function $\alpha_{1}(\cdot)$ with $\lim_{a\rightarrow\infty}\alpha_{1}(a)=0$ and a nondecreasing function $g(\cdot)$ such that the strong-mixing coefficients $\alpha(a,b)$ satisfies the inequality
@@ 4.3 Assumptions
With the data $(X(s_{1}),Y(s_{1})),\cdots,(X(s_{n}),Y(s_{n}))$, we would like to construct a prediction region for $Y(s_{n+1})$ where only the feature $X(s_{n+1})$ is known. Define $\omega_{n}(X(s_{i}))$ to be the normalized weight over samples $s_{1},\cdots,s_{n}$ so that $\sum_{i=1}^{n}\omega_{n}(X(s_{i}))=1$. The weight function can be some function that measures the similarity between $X(s)$ and $X(s_{n+1})$.
@@
$\omega_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}}),$ (5)
for all $i$; meaning that $M_{n}=\max_{1\leq i\leq n}\omega_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}})$.
The requirement assumes that the normalized weights decay at a rate faster than $n^{-\frac{1}{2}}$. As we can see, $\omega_{n}=\frac{1}{n}$ is a special case that satisfies this condition. The condition is also weaker than the requirement of $\omega_{n}=O(\frac{1}{n})$ in a related study (Xu et al., 2024). Besides, the assumption can also be inferred from that of SLSCP (Mao et al., 2024) where a GBF kernel with infinitely close data leads to uniform weights.
@@
$\begin{split}&\sum_{i=1}^{n}\|\hat{\varepsilon}(s_{i})-\varepsilon(s_{i})\|^{2}\leq\frac{\delta_{n}^{2}}{M_{n}},\\ &\|\hat{\varepsilon}(s_{n+1})-\varepsilon(s_{n+1})\|\leq\delta_{n}.\end{split}$ (6)
@@ 4.4 Coverage guarantee
Lemma 4.4 (Distance between the empirical CDF of $\varepsilon$ and $\hat{\varepsilon}$).
Under Assumption 4.1 and 4.2,
@@
Lemma 4.5 (Convergence of empirical CDF of $\varepsilon$).
Under Assumptions 4.1-4.3, with probability higher than $1-(2+M+2\sqrt{M}g(b))n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}$,
@@
where $\widetilde{F}_{n+1}(y)=\sum_{i=1}^{n}\omega_{n}(X(s_{i}))\mathbf{1}(\epsilon(s_{i})\leq y)$.
Our main theorem is the following: Using the previous lemma, we establish the asymptotic convergence of conditional coverage.
Theorem 4.6 (Conditional coverage guarantee).
Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
@@
Corollary 4.7 (Marginal Coverage).
Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
@@
Scenario 1
(a) Coverage
(b) Width
Scenario 2
(c) Coverage
(d) Width
Scenario 3
(e) Coverage
(f) Width
New Mexico
(g) Coverage
(h) Width
Georgia
(i) Coverage
(j) Width
Figure 3: The violin plots on the left show the distribution of coverage across different areas, while the plots on the right show the distribution of width. Each row represents a different scenario or location.
@@
Figure 4: The heatmaps illustrate the width of the prediction intervals for each method across the three scenarios. The width heatmap of LSCP closely matches the true residual heatmap, demonstrating its ability to capture fine details accurately.
5 Experiments
In this section, we compare our proposed LSCP method against four baseline approaches. The first method, EnbPI (Xu & Xie, 2021), is designed for time series data and applies equal weight to the most recent observations. The second method, GSCP (Mao et al., 2024), uses the entire dataset for prediction, applying equal weights to all data points. The third method, SLSCP (Mao et al., 2024), leverages k-nearest neighbors for uncertainty quantification, assigning different weights to each point based on the spatial distance between the test data and the calibration data. The fourth method, LCP (Guan, 2023), utilizes all data points for prediction, weighting them according to the similarity between the features of test data and calibration data. Both SLSCP and LCP use the Gaussian kernel as a similarity measure.
5.1 Synthetic data experiments
We begin by comparing LSCP with baseline methods across several simulated scenarios. For simplicity, we assume that the locations $s$ are uniformly sampled from the unit grid $[0,1]\times[0,1]$. The mean-zero stationary Gaussian process $X(s)$ is defined using a Matérn covariance function with variance $\sigma^{2}=1$, range $\phi=0.1$, and smoothness $\kappa=0.7$. The scenarios are as follows:
  1. 1.
    $Y(s)=X(s)+\epsilon(s)$.
  2. 2.
    $Y(s)=X(s)|\epsilon(s)|$.
  3. 3.
    $Y(s)=X(s)+\sin(\|s\|_{2})\epsilon(s)$.
These scenarios incorporate nonlinear and non-stationary settings that extend beyond the assumptions of our theoretical framework. Nevertheless, the empirical results demonstrate that LSCP consistently outperforms other methods across all scenarios.
In Table 2, LSCP achieves the target coverage of $90\%$ across all cases while maintaining significantly narrower prediction intervals. Figure 4 visualizes the interval width over the spatial domain for each scenario and the true residuals. GSCP is omitted from these plots as it produces uniform interval widths across space. A larger interval width indicates higher model uncertainty. The results show that LSCP adapts more locally, whereas the other methods exhibit smoother, more uniform patterns. Both EnbPI and SLSCP rely on $k$-nearest neighbors for constructing prediction intervals, differing primarily in their weighting schemes.
To evaluate the methods, we divide the grid into $10\times 10$ areas and compute coverage and average interval width for each, as shown in Figure 3. Both LSCP and GSCP achieve average coverage above the target $90\%$, with LSCP exhibiting the most stable and consistent distribution. LSCP also attains the smallest interval width and the highest consistency. LCP performs similarly to GSCP, as both utilize the entire calibration dataset, and as kernel bandwidth increases, LCP weights converge to GSCP’s uniform weights. Overall, LSCP produces tighter valid regions with superior consistency.
Table 2: Simulation: The table presents a comparison of the coverage and prediction interval width for five methods across three scenarios. Target coverage is $90\%$.
@@ 5.1 Synthetic data experiments
SLSCP 90.6% 1.64 88.2% 0.73 89.1% 1.07
LCP 89.6% 1.92 94.1% 0.78 91.8% 1.43
Table 3: Real data: The table presents a comparison of the coverage and prediction interval width for five methods on mobile signal data. The target coverage is $90\%$.
@@
SLSCP 89.8% 276.8 89.6% 167.8
LCP 89.7% 266.4 89.3% 166.7
5.2 Real data experiments
In the real-data experiments, we utilize mobile network measurement data from the Ookla public dataset, which records user-reported statewide mobile internet performance. Our analysis focuses on datasets from New Mexico and Georgia. The data is unevenly distributed across urban and suburban areas. We employ kernel regression as the prediction method $\hat{f}$ and compare the performance of conformal methods. The New Mexico dataset contains 24,983 observations, while the Georgia dataset includes 28,587 data points.
As shown in Table 3, our proposed LSCP method outperforms competing methods by achieving significantly narrower prediction intervals while maintaining better coverage. Similar to the synthetic experiments, we divide each state into $10\times 10$ grids and compute coverage and interval width for the test data within each grid. The violin plots in Figure 3 illustrate the grid-level distribution of these metrics, highlighting spatial performance. Unlike the table, which reports averages over all test points, the plots emphasize spatial consistency, showing higher coverage in urban areas due to denser data.
The results demonstrate that LSCP consistently achieves high coverage with narrow, uniform intervals, outperforming alternative methods. The violin plots confirm LSCP’s stability and uniformity across the spatial domain, underscoring its robustness to non-uniform data distributions. These findings align with our synthetic study, further validating LSCP’s effectiveness in diverse and complex settings.
6 Conclusion